A researcher wants to estimate the average number of days/week UNBC students engage in physical exercises in response to a national fitness magazine article suggesting that the average student in Canada does not engage in physical exercises. She is planning to select a simple random sample of students using the list of currently enrolled students at UNBC after receiving approval from the Research Ethics Board. Assuming the researcher took a sample of 50 participants and observed a sample average engagement in physical exercise as 2.5/ week, find the 99% confidence interval of this estimate if the population standard deviation is 1.5. Select one: a. 1.42 to 4.21 b. 1.95 to 3.05 c. 2.08 to 2.92 d. 2.15 to 2.85
Answer #1

We have n = 50 , xbar = 2.5 and s = 1.5

The 99% confidence interval for population mean is

xbar - E < \mu < xbar + E

Where E = t​​​​​​a/2*(s/√n)

For a = 0.01 and d.f = n -1 = 49

t​​​​​​a/2 , n -1 = t​​​​​​​​​​​0.005 , 49 = 2.68

E = 2.68*(1.5/√50) = 0.57

2.5 - 0.55 < \mu < 2.5 + 0.55

1.95 < \mu < 3.05

Option b is correct